Let a=3i-2j+5k & b=i+3j+4k. Find vector a*b & a.b
1
Expert's answer
2020-01-29T14:09:07-0500
"\bold{a\cdot b}=3\cdot 1-2\cdot 3+5\cdot 4=17"

"\bold{a\times b}=\begin{vmatrix}
 i & j & k\\
 3 & -2 & 5\\
 1 & 3 & 4\\
\end{vmatrix}"

"\bold{a\times b}=-23\bold{i}-7\bold{j}+11\bold{k}"